OpenStudy (anonymous):

Evaluate the limit:

OpenStudy (anonymous):

\[\lim_{x \rightarrow 3^{+}}\frac{ \sqrt{x ^{2}-9} }{ x-3}\]

OpenStudy (anonymous):

\[\lim_{x \rightarrow 3^{+}}\frac{ \sqrt{x ^{2}-9} }{ x-3 }*\frac{ \sqrt{x ^{2}+9} }{ \sqrt{x ^{2}+9} } = \frac{ x ^{2}-9 }{ x-3(\sqrt{x ^{2}-9)} }\] that's right?

OpenStudy (anonymous):

You should have \(\sqrt{x^2+9}\) in the denominator. Then factor the numerator and the \(x-3\) factors disappear.

OpenStudy (anonymous):

\[\lim_{x \rightarrow 3^{+}}\frac{ \sqrt{x ^{2}-9} }{ x-3 }*\frac{ \sqrt{x ^{2}+9} }{ \sqrt{x ^{2}+9} } = \frac{ x ^{2}-9 }{ x-3(\sqrt{x ^{2}-9)} }\] So true solution?

OpenStudy (anonymous):

\[\lim_{x \rightarrow 3^{+}}\frac{ \sqrt{x ^{2}-9} }{ x-3 }*\frac{ \sqrt{x ^{2}+9} }{ \sqrt{x ^{2}+9} } = \frac{ x ^{2}-9 }{ x-3(\sqrt{x ^{2}-9)} } = \frac{ (x+3)(x-3) }{x-3 (\sqrt{x ^{2}-9)} }= \frac{ (x+3) }{ \sqrt{x ^{2-9}} }\]

OpenStudy (anonymous):

Right, but it should be \[\lim_{x\to3^+}\frac{x+3}{\sqrt{x^2+9}}\] You can plug in 3 right away.

OpenStudy (anonymous):

thank you :)

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

Hold on a sec, I skimmed over your work again and I think something might be wrong with it...

OpenStudy (anonymous):

Like what?

OpenStudy (anonymous):

The main issue I'm seeing is that \[\sqrt{x^2-9}\cdot\sqrt{x^2+9}\not=x^2-9\]

OpenStudy (anonymous):

so what now ? :( I think it's true ..

OpenStudy (anonymous):

It's not, you're confusing it with \[\sqrt{x^2-9}\cdot\sqrt{x^2-9}=x^2-9\]

OpenStudy (anonymous):

Distributing, we'd get \[\sqrt{x^2-9}\cdot\sqrt{x^2+9}\not=\sqrt{x^4-81}\] which I don't think is very helpful...

OpenStudy (anonymous):

so ?

OpenStudy (anonymous):

\[\sqrt{x ^{2}-9} \times \sqrt{x ^{2}+9} = (\sqrt{x ^{2}-9)^{2}} = (x ^{2}-9)\]

OpenStudy (anonymous):

Because the square root will go with quadrature

OpenStudy (anonymous):

You're saying that \[(x^2-9)(x^2+9)=(x^2-9)^2\] which is not the case. Instead of rewriting the expression algebraically to evaluate the limit, you could instead plug in numbers increasingly closer to \(x=3\), coming from the right. In other words, evaluate the expression for \(x=3.1,~3.01,~3.001\) and so on.

OpenStudy (anonymous):

this is my solution

OpenStudy (anonymous):

The problem still lies in the step between the first two lines. The algebra doesn't work out like that. Also, the limit is not finite.

OpenStudy (anonymous):

Could it be that the problem of the question may be what is wrong in it?

OpenStudy (anonymous):

I'm saying that there's no way to algebraically manipulate the expression to allow you to remove the discontinuity at \(x=3\).

OpenStudy (anonymous):

limit does not exist (or +infinity in this case) since x -> 3+, (x-3) is always positive so we can do this sqrt(x^2-9) / sqrt( (x-3)^2 ) and x^2-9 is also positive, so we can do this sqrt[ (x^2-9) / (x-3)^2 ] sqrt[ (x+3) / (x-3) ] when evaluating the limit, it's positive infinity

OpenStudy (jdoe0001):

hmmm yes, in the end you'd need to evaluate by checking a few values for "x"

OpenStudy (jdoe0001):

\(\bf \large{\cfrac{\sqrt{x^2-3^2}}{x-3}\implies \cfrac{\sqrt{(x-3)(x+3)}}{x-3}\implies \cfrac{(x-3)^{\frac{1}{2}}(x+3)^{\frac{1}{2}}}{(x-3)^1} \\ \quad \\ (x-3)^{\frac{1}{2}}\cdot (x-3)^{-1}(x+3)^{\frac{1}{2}}\implies (x-3)^{\frac{1}{2}-1}(x+3)^{\frac{1}{2}} \\ \quad \\ (x-3)^{-\frac{1}{2}}(x+3)^{\frac{1}{2}}\implies \cfrac{\sqrt{x+3}}{\sqrt{x-3}}}\) but pretty much what sourwing said